# Curious Causality-Seeking Agents Learn Meta Causal World

### v1

@@ 1 Introduction
To identify the underlying world rules, traditional causal models rely predominantly on observational data (pmlr-v216-lippe23a), implicitly assuming a fixed causal structure. In open-ended environments, although the ground-truth causal laws remain unchanged, the observational data is collected through local, context-limited exploration. Under conditions of partial observability, this invariance becomes obscured within a limited observational window, giving rise to the illusion of shifting causal relationships—a rule valid in one context may break down in another. For instance, the "push $\rightarrow$ open" relation holds for an unlocked door, but this relation vanishes when the door is locked. In this case, the same action produces no effect, leaving the true causal mechanism ambiguous. Traditional causal modeling in such scenarios faces two main challenges: (1) methods that learn a single uniform (time series-based) causal graph neglect these context-dependent changes (NEURIPS2023_65496a49), and (2) lack of an active curiosity-driven exploration strategy, which makes it struggle to adapt to discover the global causal rules in rich, unseen variations of potential open-ended worlds (huang2020causal; hwang2024fine).
To deal with these two challenges, we propose the Meta-Causal Graph to describe the world, which is a unified structure that captures how causal relationships evolve across different states. Rather than relying on a uniform graph, the Meta-Causal Graph is a minimal representation that contains multiple causal subgraphs, each corresponding to a particular meta state. Transitions between these meta states “trigger” the activation of the appropriate subgraph, allowing us to model shifts in causal influence, such as when a locked door severs the "push $\rightarrow$ open" link. We introduce a framework termed Curious Causality-Seeking Agent, which actively explores interventions to uncover the Meta-Causal Graph. The main contributions of this work are given below:
@@ 2 Preliminaries
Let $D=([p],E)$ be a DAG with vertex set $[p]$ and edge set $E$, and $I\subset[p]$ an intervention target. The intervention graph of $D$ is the DAG $D^{(I)}=([p],E^{(I)})$, where $E^{(I)}:=$$\{a\to b\mid a\to b\in E,b\notin I\}.$
@@ 3.1 Meta-Causal Graph
    Causality skeleton matrices of causal subgraphs are sufficiently different. For all $u\neq u^{\prime}$, $M_{u}\neq M_{u^{\prime}}$, i.e., there exists an index $(i,j)$ such that $M_{u}[i,j]\neq M_{u^{\prime}}[i,j]$.
@@ 4.1 Meta-Causal Graph Learning from Experience
We learn causal subgraphs under latent meta states from agent experience. Specifically, the process contains two steps: (1) identifying the latent meta state, which activates a specific causal subgraph, and (2) learning the subgraph from the agent’s explored experience data.
@@
where $\text{sg}(\cdot)$ is the stop-gradient operator. The first term is the reconstruction loss and the second term is the commitment loss to avoid the output of the encoder growing arbitrarily (van2017neural).
@@ 4.3 Transition Probability Learning
The previous subsections describe how to identify the parent set of each state variable, $Pa_{\mathcal{G}_{u}}(X_{j})$, using the learned causal skeleton matrix $M_{\theta}$. In this section, we demonstrate how to leverage the confirmed causal skeleton to quantify inter-variable causal effects, thereby grounding and validating the transition probabilities of the world model for next-state prediction. The model is optimized via a maximum likelihood estimation objective:
@@ 6.1 Experiment Setup
We compare our method against several baselines: MLP, a dense model that predicts transition dynamics $p(x_{t+1}|x_{t},a_{t})$; Modular, a modular network with separate modules for each state variable; GNN (kipf2019contrastive), a graph neural network for predicting transition dynamics; NPS (alias2021neural), which learns sparse and modular dynamics; CDL (wang2022causal), which learns a static causal model from data; GRADER (ding2022generalizing), which learns a static causal model using conditional independence tests; NCD (hwang2024fine), a neural causal discovery algorithm for learning causal graphs for each sample; and FCDL learns causal graph to facilitate the robustness for reinforcement learning. Please check Appendix B for details of the experimental setup.
@@ A.3 Proof of Theorem 1
Moreover, since $M_{u_{1}}\neq M_{u_{2}}$, there exists at least one position $(i,j)$ where $M_{u_{1}}[i,j]\neq M_{u_{2}}[i,j]$.
@@ A.6 Disscussion on Assumption 1
where $\mathbb{I}[\cdot]$ is the indicator function, ensuring that an edge $(i,j)$ exists in the learned graph if and only if it exists in at least one contributing subgraph.
@@ A.8 Reachability of Interventions
Reachability Analysis: Starting from $\mathbf{z}_{0}=[1,0,0,0]^{\top}$ (state $(0,0)$):

### v2

@@ 1 Introduction
To identify the underlying world rules, traditional causal models rely predominantly on observational data (pmlr-v216-lippe23a), implicitly assuming a fixed causal structure. In open-ended environments, although the ground-truth causal laws remain unchanged, the observational data is collected through local, context-limited exploration. Under conditions of partial observability, this invariance becomes obscured within a limited observational window, giving rise to the illusion of shifting causal relationships—a rule valid in one context may break down in another. For instance, the "push → open" relation holds for an unlocked door, but this relation vanishes when the door is locked. In this case, the same action produces no effect, leaving the true causal mechanism ambiguous. Traditional causal modeling in such scenarios faces two main challenges: (1) methods that learn a single uniform (time series-based) causal graph neglect these context-dependent changes (NEURIPS2023_65496a49), and (2) lack of an active curiosity-driven exploration strategy, which makes it struggle to adapt to discover the global causal rules in rich, unseen variations of potential open-ended worlds (huang2020causal; hwang2024fine).
To deal with these two challenges, we propose the Meta-Causal Graph to describe the world, which is a unified structure that captures how causal relationships evolve across different states. Rather than relying on a uniform graph, the Meta-Causal Graph is a minimal representation that contains multiple causal subgraphs, each corresponding to a particular meta state. Transitions between these meta states “trigger” the activation of the appropriate subgraph, allowing us to model shifts in causal influence, such as when a locked door severs the "push → open" link. We introduce a framework termed Curious Causality-Seeking Agent, which actively explores interventions to uncover the Meta-Causal Graph. The main contributions of this work are given below:
@@ 2 Preliminaries
[…]
@@ 3.1 Meta-Causal Graph
[…]
@@ 4.1 Meta-Causal Graph Learning from Experience
We learn causal subgraphs under latent meta states from agent experience. Following FCDL’s pioneering use of the VQ-VAE architecture for causal-graph discovery (hwang2024fine), we adopt the original VQ-VAE framework of van den Oord et al. (van2017neural) to learn the causal subgraph for each meta-state. Specifically, the process contains two steps: (1) identifying the latent meta state, which activates a specific causal subgraph, and (2) learning the subgraph from the agent’s explored experience data.
@@
[…]
@@ 4.3 Transition Probability Learning
The previous subsections describe how to identify the parent set of each state variable, $Pa_{\mathcal{G}_{u}}(X_{j})$, using the learned causal skeleton matrix $M_{\theta}$. In this section, we demonstrate how to leverage the confirmed causal skeleton to quantify inter-variable causal effects, thereby grounding and validating the transition probabilities of the world model for next-state prediction. We train our model by maximizing the log-likelihood of the observed sequences, thereby learning the transition dynamics.
@@ 6.1 Experiment Setup
We compare our method against several baselines: MLP, a dense model that predicts transition dynamics $p(x_{t+1}|x_{t},a_{t})$; Modular, a modular network with separate modules for each state variable; GNN (kipf2019contrastive), a graph neural network for predicting transition dynamics; NPS (alias2021neural), which learns sparse and modular dynamics; CDL (wang2022causal), which learns a static causal model from data; GRADER (ding2022generalizing), which learns a static causal model using conditional independence tests; NCD (hwang2024fine), a neural causal discovery algorithm for learning causal graphs for each sample; and FCDL (hwang2024fine) learns causal graph to facilitate the robustness for reinforcement learning. Please check Appendix B for details of the experimental setup.
@@ A.3 Proof of Theorem 1
[…]
@@ A.6 Disscussion on Assumption 1
[…]
@@ A.8 Reachability of Interventions
Reachability Analysis: Starting from $\mathbf{z}_{0}=[1,0,0,0]^{\top}$ (state (0,0)(0,0)( 0 , 0 )):